What is the force on na electron in an electric field of 200 N/C?


1
Expert's answer
2022-03-13T18:47:21-0400

Answer

the force on an electron in an electric field of 200 N/C

"F=qE\\=1.6*10^{-19}*200\\=320*10^{-19}N"